OpenStudy (alina123):

7th grade baseline the temperature decreases a total of 26.4 degrees celsius over 2.2 days in aruba in the middle of the winter. write a number that represents how much the temperature changes each hour

OpenStudy (texaschic101):

2.2 days is how many hrs ? That is where I am stuck. It is a little over 48 hrs.... or 1 hr = .04166 rounds to .0417 days I am trying to make a proportion.. 26.4/2.2 = x/.0417 I am not even sure if I set this up right.....you will need someone who knows more about what they are doing then me...sorry

OpenStudy (texaschic101):

or do you just divide ??

OpenStudy (johnweldon1993):

Well if we are breaking 1 day into a fraction...lets do 24 hours / 5 = 4.8 hours That's 20% of 1 day... So in 2 days there are 48 + 4.8 = 52.8 hours.... So I would just do \[\large \frac{26.4}{52.8} = \frac{x}{1}\] \[\large x = \frac{26.4}{52.8}\]

OpenStudy (johnweldon1993):

Sorry...up there I meant *in 2.2 days there are 48 + 4.8 = 52.8 hours

OpenStudy (texaschic101):

oh...you changed your days to hrs instead of the way I set it up. Your way makes more sense...lol

OpenStudy (alina123):

So per hour what would that be

OpenStudy (johnweldon1993):

Well when you solve that for 'x' what do you get?

OpenStudy (alina123):

Oh never mind I read the proportion

OpenStudy (texaschic101):

just do the division.. 26.4 / 52.8

OpenStudy (alina123):

Ok 0.5 rounded

OpenStudy (johnweldon1993):

Right... So that equates to .5 degrees Celsius an hour That would mean .5 * 52.8 would be the 26.4 And that checks out

OpenStudy (texaschic101):

glad you were here johnweldon :)

OpenStudy (alina123):

Ok thanks a bunch

OpenStudy (johnweldon1993):

Just a tag away :P lol